OpenStudy (anonymous):

I nedd helpppp!!!! The vertices of triangle ABC on a coordinate grid are A(-1, 1), B(-6,6), and C(-1,6). The vertices of triangle AQR are A(-1,1), Q(-4,4), and R(-1,4). Which proportion can be used as a step in proving that triangle ABC is similar to triangle AQR?

OpenStudy (anonymous):

AR/CR = 5/3√2 AC/AQ= 5/3√2 AB/AQ=5/3 BC/QR=5/2

OpenStudy (anonymous):

@tcarroll010